THE MICROWAVE SPECTRUM AND MOLECULAR STRUCTURE OF TRIFLUOROETHYLENE-HYDROGEN FLUORIDE

\underline{MARK D. MARSHALL}, HELEN O. LEUNG, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

The rotational spectra of four isotopomers of the 1,1,2–trifluoroethylene-hydrogen fluoride complex (HF-CHFCF_{2}, DF-CHFCF_{2}, HF-^{13}CHFCF_{2}, and HF-CHF^{13}CF_{2}) have been collected in the 7–22 GHz region with a Fourier transform microwave spectrometer. The complexes are generated using pulsed jet expansion of a mixture of HF (or DF), CHFCF_{2}, and Ar, and the two ^{13}C-containing isotopomers are observed in natural abundance. Only a-type transitions are seen for all isotopomers, and for most lines in the DF-containing isotopomer, deuterium nuclear quadrupole hyperfine structure is not resolved within the experimental linewidth of 10–15 kHz. Both the Kraitchman substitution coordinates and the rotational constants are consistent with a planar structure in which a hydrogen bond is formed between H in HF and the F on C-2 in CHFCF_{2} with a secondary interaction between the H (geminal to the F involved in the hydrogen bond) in CHFCF_{2} and the F in HF.